OpenStudy (anonymous):

(5x-3)(4x+4)=20x^2+8x-12 is this true or false

OpenStudy (angela210793):

true...but...why r u asking such q? O.o

OpenStudy (anonymous):

for lovelesstime check this

OpenStudy (anonymous):

I said 9x + 1 and I still say it is! lol I don't understand where you got the 2 . Maybe my math isn't as advanced as that but I still say 9x+1 off of my level of education

OpenStudy (anonymous):

its xpower2 and not 2

OpenStudy (angela210793):

where we got 2? wht 2? O.o

OpenStudy (angela210793):

\[20x^2+8x-12\]

OpenStudy (anonymous):

I don't know where you get xpower 2 or why you multiple them

OpenStudy (angela210793):

cause that's wht u have to do *multiply* and 5x*4x=20*x*x=20x^2

OpenStudy (anonymous):

angela210793 thanks

OpenStudy (angela210793):

ur welcome :P :)

OpenStudy (anonymous):

So then why isn't it (5x-3)(4x+4) = 5x - 3 * 4x + 4 if you multiply.

OpenStudy (anonymous):

you removed the brackets

OpenStudy (anonymous):

I'm seeing it as the the distributive property with an invisible 1.

OpenStudy (angela210793):

see...(a-b)*(c+d)=ac+ad-bc-cd get it

OpenStudy (anonymous):

No. i'm not that advanced yet.

OpenStudy (anonymous):

then you must not say others wrong unless you know it perfectly right

OpenStudy (angela210793):

U don't have to be advanced for this...all u have to do is multiply a with c and then a with d after ur done with a u start with b,multiply b with c and then b with d Got it now?

OpenStudy (anonymous):

But when I at my level look at that I see 9x + 1 it's not wrong nor right I only said you were wrong because you said I was wrong. And no Angela I don't get it now.

OpenStudy (anonymous):

then better leave that

OpenStudy (angela210793):

let me search for a link..maybe u'll get it

OpenStudy (anonymous):

This is stupid it's just a math question. Angela I'm 14 not even in highschool yet I will in August do not make me learn in the summer.

OpenStudy (angela210793):

C'mon..there will be no harm if u learn this....u'll need it for all ur life http://www.mathsisfun.com/algebra/polynomials-multiplying.html take a look at this

OpenStudy (anonymous):

I have not learned it yet, I do not care I am not clicking the link.

OpenStudy (anonymous):

angela210793 that makes you great .for giving good information

OpenStudy (angela210793):

Ok....a thank you for helping would be enough..... Sorry for trying to help you X(